an arc of a circle of radius 7cm subtend on angle of 120 degree at the centre, find its perimeter
Answer by Seutip(231)

r = 7cm
ϴ=120°=(120°)(2π/360)=2π/3
Do you mean the length of the arc as the perimeter?

s (length of the arc) = ϴ*r
s=rϴ
s=(7) (2π/3)
s=14π/3=14.66cm
Hope that helps!